A font sampler

Alan Jeffrey

v0.10

1 Introduction

This document describes the font test document fontsmpl.tex and the accompanying package fontsmpl.sty. This produces a test of a font family, printing a sample text, a table of accents, and a sample of commands such as \pounds.

It can be used in two ways. The fontsmpl package provides the command \fontsample, which produces a sample of the current font.

The fontsmpl document prompts interactively for a font family (for example ‘cmr’) and produces a sample of that family.